Circular No. 6891

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1806–246 F. E. Marshall and T. Strohmayer, Goddard Space Flight Center;

and R. Remillard, Massachusetts Institute of Technology, on behalf of the RXTE Science Operations Center, report the detection with the Rosssi X-ray Timing Explorer of a bright, soft x-ray transient (designated XTE J1806–246) at a position consistent with that of the SAS source 2S 1803–245 and a possibly related source of isolated bursts, SAX J1806.8–2435 (IAUC 6867): “The 2–12-keV flux measured with the All-Sky Monitor increased as follows: Apr. 16.66 UT, 31 mCrab; 18.27, 60; 22.42, 131; 24.45, 356; 28.93,

417. The 3–20-keV spectrum measured with the Proportional Counter Array (PCA) on Apr. 28.81 is well described as thermal bremsstrahlung emission with kT = 4.3 keV. The source position determined with the PCA is α = 18^{h}06^{m}51^{s}, δ = −24^{o}35′.1 (equinox 2000.0) with an estimated uncertainty of 1′.5. Observations at other wavelengths are encouraged.”

SUPERNOVA 1998bp IN NGC 6495 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, communicate: “A spectrum (range 360–750 nm) of SN 1998bp was taken by M. Calkins with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on Apr. 30.4 UT. The spectrum confirms this object (cf. IAUC 6890) to be a type-Ia supernova near maximum light. In addition to a strong Si II 615-nm absorption, there is a very deep Si II 580-nm feature, typical of subluminous, fast-declining type-Ia supernovae. The Si line-depth ratio [R(Si); Nugent et al. 1995, Ap.J. 455, L147] is 0.62, similar to that of SN 1991bg near maximum. However, the spectrum does not resemble that of SN 1991bg in the blue; it is more similar to that of SN 1992A. The catalogue recession velocity for NGC 6945 is 3127 km/s (Di Nella et al. 1995, A.Ap. Suppl. 113, 151), which yields an expansion velocity from the most prominent Si II feature (rest 635.5 nm) of −10 400 km/s.”

SUPERNOVAE Unfiltered CCD measurements by K. Sárneczky and L. Kiss, Szeged,

Hungary, with a 0.90-m Schmidt telescope (comparison-star R magnitudes from USNO A1.0 catalogue): SN 1998V in NGC 6627, Apr. 4.05 UT, 16.2; SN 1998ab in NGC 4704, Apr. 5.11, 15.9.

1998 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green